# Physics-Informed Wireless Imaging with Implicit Neural Representation in RIS-Aided ISAC System

### v1

@@ Abstract
Wireless imaging is emerging as a key capability in next-generation integrated sensing and communication (ISAC) systems, supporting diverse context-aware applications. However, conventional imaging approaches, whether based on physical models or data-driven learning, face challenges such as accurate multipath separation and representative dataset acquisition. To address these issues, this study explores the use of implicit neural representation (INR), a paradigm that has achieved notable advancements in computer vision, for wireless imaging in reconfigurable intelligent surface-aided ISAC systems. The neural network of INR is specifically designed with positional encoding and sine activation functions. Leveraging physics-informed loss functions, INR is optimized through deep learning to represent continuous target shapes and scattering profiles, enabling resolution-agnostic imaging with strong generalization capability. Extensive simulations demonstrate that the proposed INR-based method achieves significant improvements over state-of-the-art techniques and further reveals the focal length characteristics of the imaging system.
@@ I Introduction
Integrated sensing and communication (ISAC) is one of the primary scenarios for future communication systems [1]. Rather than being confined to locating point targets, ISAC has shown a tendency to sense fine-grained shape and scattering characteristics of extended targets through wireless imaging [2, 3]. Additionally, reconfigurable intelligent surfaces (RISs) have become a key enabler for ISAC to adjust electromagnetic fields, providing diverse sensing perspectives and channel state information (CSI) measurements [4]. This study aims to realize extended target imaging within RIS-aided ISAC systems.
@@
Wireless imaging has been widely studied in literature, where image domain optimization problems are formulated and solved to obtain imaging results. Traditional model-based imaging methods include Fourier transform (FT) and compressed sensing (CS), exhibiting distinct advantages in terms of computational complexity and imaging accuracy. However, both approaches require multipath extraction to retain only the transmitter (TX)-ROI-RIS-receiver (RX) path [5], potentially losing useful information for image generation. Moreover, channel modeling errors may significantly distort the imaging results [6, 3].
@@
In contrast, data-based methods reduce reliance on physical models by learning the relationship between CSI measurements and target images from a large dataset [7]. Although deep learning (DL) techniques have achieved improved imaging performance, constructing such a large dataset is challenging since ground-truth images are difficult to obtain. Furthermore, transferring well-trained model parameters to new imaging scenarios, targets, and resolutions is complicated [8]. While researchers have attempted to integrate physical models into learning [9], these challenges remain inadequately addressed.
@@
Recently, implicit neural representation (INR) has shown excellent performance in computer vision field [10]. Specifically, the term implicit neural representation can be decomposed into: 1) implicit representation, indicating that INR uses continuous functions to represent sensing results rather than explicit and discretized forms such as images and point clouds, and 2) neural representation, meaning that these implicit functions are represented by neural networks (NNs).
In this study, we introduce INR for wireless imaging in RIS-aided ISAC systems, transforming traditional image domain optimization into the optimization of NN parameters, which can be performed using DL techniques under the supervision of differentiable physical models. INR alleviates the need for multipath extraction and does not require ground-truth images or large training datasets, making it significantly superior to traditional model-based and data-based methods. By employing INR, the continuous target shape and scattering characteristics can be learned through NN training, enabling INR to render images of arbitrary resolutions, which is infeasible for state-of-the-art imaging techniques.
@@ II-A System Configuration
We consider a RIS-aided ISAC system operating in a 3D space $x,y,z]^{\text{T}}\in\mathbb{R}^{3}$, as illustrated in Fig. 1. A full-duplex BS is employed to enable both user communication and target imaging. The BS’s TX and RX are equipped with uniform linear arrays consisting of $N_{\text{t}}$ and $N_{\text{r}}$ antennas, respectively, with half-wavelength spacing $\lambda/2$, where $\lambda$ is the center subcarrier wavelength. Self-interference at the BS is assumed to be mitigated by isolated TX and RX arrays [11]. The RIS is composed of $N_{\text{s}}$ elements, each with a size of $\xi_{\text{s}}\times\xi_{\text{s}}$. Its phase shifts are $\bm{\omega}=[\omega_{1},\omega_{2},\ldots,\omega_{N_{\text{s}}}]^{\text{T}}\in\mathbb{C}^{N_{\text{s}}\times 1}$. The positions of the TX, RX, and RIS are assumed to be known, whereas other environmental components are unknown but considered static.
@@
The region of interest (ROI) is a predefined area where potential targets may be located. It can be divided into $N_{\text{v}}$ voxels, whose side length is $\xi_{\text{v}}$, which is much smaller than $\lambda/2$ [12, 13]. The ROI image is represented as $\bm{\sigma}=[\sigma_{1},\sigma_{2},\ldots,\sigma_{N_{\text{v}}}]^{\text{T}}\in\mathbb{R}^{N_{\text{v}}\times 1}$, where $\sigma_{n_{\text{v}}}$ denotes the scattering coefficient of the $n_{\text{v}}$-th voxel, and $\sigma_{n_{\text{v}}}=0$ if the voxel contains no object. Communication and imaging functions operate in a time-division manner, and $N_{\text{s}}$ symbol intervals are used to generate ROI images.
@@
Figure 1: The proposed RIS-aided ISAC system for wireless imaging.
II-B Signal and Channel Models
@@
The BS performs target sensing by transmitting signals from the TX and receiving the echoes at the RX. We assume that RIS phases are altered using $K$ different configurations for imaging, given by $\bm{\Omega}=[\bm{\omega}_{1},\bm{\omega}_{2},\ldots,\bm{\omega}_{K}]^{\text{T}}\in\mathbb{C}^{K\times N_{\text{s}}}$. While RIS phase optimization for imaging has been studied in [13], this work focuses on imaging algorithm design and assumes that $\bm{\Omega}$ is randomly generated in advance. When employing the $k$-th RIS configuration $\bm{\omega}_{k}$ and transmitting the $n_{\text{t}}$-th signal $\mathbf{x}_{n_{\text{t}}}$, the received signal at the RX is expressed as
@@
where $g_{\text{sen}}={\lambda\sqrt{G_{\text{sen}}}}/{\sqrt{4\pi}}$, and $G_{\text{sen}}$ is the combined antenna gain of the TX and the RX. $\mathbf{H}_{\text{tx-roi}}$ and $\mathbf{H}_{\text{tx-ris}}$ denote the channels from the TX to the ROI and the RIS, respectively. $\mathbf{H}_{\text{roi-rx}}$ and $\mathbf{H}_{\text{ris-rx}}$ are the channels from the ROI and the RIS to the RX, respectively. The terms $\mathbf{H}_{\text{tx-roi-ris-rx},k}$ and $\mathbf{H}_{\text{tx-ris-roi-rx},k}$ represent twice-bounce paths, given by
@@
$\mathbf{H}_{\text{tx-ris-roi-rx},k}\!$ $=g_{\text{sen}}\mathbf{H}_{\text{tx-ris}}\text{diag}(\widetilde{\bm{\omega}}_{k})\mathbf{H}_{\text{ris-roi}}\text{diag}(\bm{\sigma})\mathbf{H}_{\text{roi-rx}},$ (4b)
where $\mathbf{H}_{\text{roi-ris}}=\mathbf{H}_{\text{ris-roi}}^{\text{T}}$ is the channel between the RIS and the ROI. ${\mathbf{H}}_{\text{others},k}$ includes multipaths involving more than two bounces or scatterers beyond the ROI and RIS, representing the unknown portion of the sensing channel ${\mathbf{H}}_{\text{sen},k}$. The next section aims to retrieve $\bm{\sigma}$ from the measurements of ${\mathbf{H}}_{\text{sen},k}$.

### v2

# Physics-Informed Implicit Neural Representation for Wireless Imaging in RIS-Aided ISAC System

@@ Abstract
Wireless imaging has become a vital function in future integrated sensing and communication (ISAC) systems. However, traditional model-based and data-driven deep learning imaging methods face challenges related to multipath extraction, dataset acquisition, and multi-scenario adaptation. To overcome these limitations, this study innovatively combines implicit neural representation (INR) with explicit physical models to realize wireless imaging in reconfigurable intelligent surface (RIS)-aided ISAC systems. INR employs neural networks (NNs) to project physical locations to voxel values, which is indirectly supervised by measurements of channel state information with physics-informed loss functions. The continuous shape and scattering characteristics of targets are embedded into NN parameters through training, enabling arbitrary image resolutions and off-grid voxel value prediction. Additionally, three issues related to INR-based imager are further addressed. First, INR is generalized to enable efficient imaging under multipath interference by jointly learning image and multipath information. Second, the imaging speed and accuracy for dynamic targets are enhanced by embedding prior image information. Third, imaging results are employed to assist in RIS phase design for improved communication performance. Extensive simulations demonstrate that the proposed INR-based imager significantly outperforms traditional model-based methods with super-resolution abilities, and the focal length characteristics of the imaging system is revealed. Moreover, communication performance can benefit from the imaging results. Part of the source code for this paper can be accessed at https://github.com/kiwi1944/INRImager
@@ I Introduction
Integrated sensing and communication (ISAC) has been incorporated as one of the primary application scenarios for sixth-generation communication systems [1, 2]. Various sensing functions, including recognition, localization, and mapping, are anticipated to be integrated into future communication systems [3, 4, 5]. Rather than being confined to detecting and locating point targets, ISAC has shown a tendency toward sensing fine-grained shape and scattering characteristics of extended targets through wireless imaging [6, 7, 8]. It offers notable benefits in all-day sensing and privacy protection without the use of visible light, inspiring numerous applications such as health monitoring and low-altitude surveillance [9, 10].
Wireless imaging in traditional communication systems has been hindered by the need for large antenna arrays and short imaging distances [11]. Recently, reconfigurable intelligent surfaces (RISs) have been recognized as a key enabler for ISAC to partially address these issues by employing large, low-cost, and energy-efficient arrays in the vast communication environment [12, 13, 14]. Additionally, RIS phase variations can manually adjust the electromagnetic field to provide diverse sensing perspectives and channel state information (CSI) measurements [15]. Given these advantages, we aim to realize extended target imaging within RIS-aided ISAC systems.
Figure 1: Relationship between model-based, data-driven DL, and INR methods.
Wireless imaging extracts target information from CSI measurements. Generally, the measurement model can be expressed as
$\mathbf{y}=f(\bm{\sigma})+\mathbf{n},$ (1)
where $\mathbf{y}\in\mathbb{C}^{N_{\text{m}}}$ is the CSI measurements, and $N_{\text{m}}$ denotes the measurement number. The scattering coefficient image of the region of interest (ROI) is denoted as $\bm{\sigma}\in\mathbb{R}^{N_{\text{v}}}$, where $N_{\text{v}}$ is the number of voxels. $f(\cdot)$ is the forward model describing the relationship between $\bm{\sigma}$ and CSI, which varies with system configurations. $\mathbf{n}\in\mathbb{C}^{N_{\text{m}}}$ represents measurement noise. Various imaging techniques aim to solve the inverse problem of (1) by optimizing the image $\bm{\sigma}$, formulated as
$\text{(P1)}\quad\bm{\sigma}^{\star}=\mathop{\arg\!\min}\limits_{\bm{\sigma}}~\mu(f(\bm{\sigma}),\mathbf{y})+\rho(\bm{\sigma}),$
where $\mu(f(\bm{\sigma}),\mathbf{y})$ measures the error between $f(\bm{\sigma})$ and $\mathbf{y}$, and $\rho(\bm{\sigma})$ is the regularization term capturing the image prior.
@@
In traditional model-based imaging methods, Fourier transform (FT) and compressed sensing (CS) both require the knowledge of the forward model $f(\cdot)$. Specifically, FT implements $f(\cdot)$ as a spatial integral, whereas $f(\cdot)$ should be discretized into a high-dimensional sensing matrix in CS theory [16, 17]. According to [18], FT- and CS-based methods exhibit distinct advantages in terms of computational complexity, imaging accuracy, and measurement requirements. However, both approaches require multipath extraction to preserve only the transmitter (TX)-ROI-RIS-receiver (RX) path [18, 15, 9], potentially discarding target information involved in other multipath components. Moreover, errors in modeling $f(\cdot)$ may significantly distort the imaging results [16, 10].
@@
In contrast, data-driven deep learning (DL) methods reduce reliance on $f(\cdot)$ by learning the relationship between $\mathbf{y}$ and $\bm{\sigma}$ from a large dataset containing numerous $(\mathbf{y},\bm{\sigma})$ pairs [19]. Although DL techniques have demonstrated superior imaging performance, constructing such a large dataset is challenging since the ground-truth (GT) ROI image is difficult to obtain [20]. Furthermore, transferring well-trained model parameters to new imaging environments, targets, and resolutions is complicated [21], requesting retraining the NNs. This has limited the application of data-driven DL methods. While researchers have attempted to integrate physical models into learning [22, 23], the above challenges in model-based and data-driven DL methods remain inadequately addressed.
@@
Recently, implicit neural representation (INR) has shown excellent performance in optical 3D reconstruction [24] and has been applied to communication systems for radio map generation [25], data suppression [26], and RIS phase design [27]. INR leverages a neural network (NN) to learn a continuous function that maps input coordinates to their corresponding signal values, thereby encoding the signal or scene within NN parameters. Specifically, the term implicit neural representation can be decomposed into: 1) implicit representation, indicating that INR uses implicit and continuous functions to represent sensing results rather than explicit and discretized forms such as images and point clouds, and 2) neural representation, meaning that these implicit functions are realized by NNs.
Inspired by the great success of INR in the above fields, this study extends its application to wireless imaging by implicitly extracting and representing radio images using NNs. Consequently, the image optimization problem (P1) is transformed into the optimization of NN parameters, which can be executed using DL techniques and accelerated by common GPU resources in future communication systems [28].
Compared with model-based methods, INR supports arbitrary differentiable forward models and mitigates the need for multipath extraction. Although DL techniques are employed, INR does not require GT images or large training datasets [19], since it is indirectly supervised by CSI measurements $\mathbf{y}$ using the physical model $f(\cdot)$ [29]. By employing INR, the continuous shape and scattering characteristics of targets can be learned through NN training. This information is implicitly embedded into NN parameters, allowing INR to generate images of arbitrary resolutions, which is infeasible for traditional methods. This feature advances the development of super-resolution and off-grid imaging [30]. The conceptual relationship between model-based, data-driven DL, and INR-based methods is illustrated in Fig. 1.
Despite these advantages, several additional issues should be addressed:
First, INR-based imagers still face the challenge of inaccurate forward models, especially in complex environments where multipath components, which are difficult to model in $f(\cdot)$, introduce significant interference. Although such interference can be suppressed through multipath extraction [18, 17], the estimation accuracy is not guaranteed. Special TX/RX placements and transmission directions are utilized in [9], but they may be impractical in communication systems. Complex algorithms are employed to design ISAC beamformers in [6, 31], but the effects of RISs are not considered. Moreover, NNs have been used to approximate $f(\cdot)$ [20, 32], requiring large training datasets, which counteract the benefits of INR. Alternatively, our study proposes to model multipath interference as trainable parameters within INR, enabling the joint learning of the ROI image and multipath components.
Second, since INR encodes image information into NN parameters, it requires online training for each target, leading to delays in producing imaging results, which is critical for dynamic target imaging. Although hierarchical sampling [24] and multiresolution hash encoding [33] have been developed to accelerate NN training for INR, they significantly increase algorithmic complexity. Unlike the location variation of point targets, the dynamics of extended targets typically manifest as changes in shape, suggesting potential temporal correlations across adjacent time frames [34, 17]. Exploiting these correlations, [35] has achieved high-accuracy medical imaging and image variation capture. Inspired by this, we propose to utilize prior image information captured in previous instants to accelerate dynamic target imaging in subsequent steps.
Third, while previous studies have focused on enhanced imaging within communication systems [18, 17, 36], the benefits of imaging for communication remain underexplored. [37] proposes using wireless depth images to detect user locations and optimize RIS radiation beams. Similarly, human body images are reconstructed and input to an NN to detect the desired sensing regions in [38]. According to [39], the knowledge of the surrounding environment, which can be supplemented by INR-based imaging results, is crucial for communication-oriented RIS phase design, leading to improved spectral efficiency (SE). In this study, we respectively address the mechanisms of imaging-augmented communication by leveraging the sensing results for users inside and outside the ROI.
To summarize, our key contributions are as follows: We pioneer the application of INR to wireless imaging in RIS-aided ISAC systems, employing DL methods under the supervision of physical models. Free of dataset construction, image information is directly extracted from CSI measurements and embedded into NN parameters through training, generating high-accuracy images at arbitrary resolutions, which traditional methods cannot achieve. In addition, the following issues are also addressed:
  • Imaging under Multipath Interference: We first eliminate ROI-irrelevant multipath components through background calibration. Then, the INR-based imager is generalized to model ROI-related interference as trainable parameters, which is learned during the imaging process.
  • Successive Imaging for Dynamic Targets: We embed prior image information from previous time instants into subsequent imaging through INR parameter initialization, thereby accelerating NN training, reducing output delay, and enhancing imaging performance.
  • Imaging-Augmented Communication: We leverage INR-based imaging results to complementary environmental information, which supports RIS phase design to enhance the communication performance for users inside or outside the ROI.
The remainder of this paper is structured as follows: Sec. II introduces the RIS-aided ISAC system, and Sec. III presents INR-based wireless imaging method. Sec. IV addresses the issues of multipath interference and dynamic target imaging. Sec. V explores imaging-augmented communication performance. Simulation results and conclusions are provided in Sec. VI and Sec. VII, respectively.
Notations—Scalars (e.g., $a$) are denoted in italics, vectors (e.g., $\mathbf{a}$) in bold, and matrices (e.g., $\mathbf{A}$) in bold capital letters. The $\ell_{1}$- and $\ell_{2}$-norms of $\mathbf{a}$ are denoted by $\|\mathbf{a}\|_{1}$ and $\|\mathbf{a}\|_{2}$, respectively. $j=\sqrt{-1}$ is the imaginary unit. $\text{diag}{(\mathbf{a})}$ constructs a diagonal matrix from the elements of $\mathbf{a}$. The transpose and Hermitian operators are $(\cdot)^{\text{T}}$ and $(\cdot)^{\text{H}}$, respectively.
@@ II System Model
We consider a RIS-aided ISAC system operating in a 3D space $\mathbb{R}^{3}=\{[x,y,z]^{\text{T}}:x,y,z\in\mathbb{R}\}$, as illustrated in Fig. 2. A full-duplex BS is employed to enable both user communication and target imaging, using orthogonal frequency division multiplexing (OFDM) signals. The BS’s TX and RX are equipped with uniform linear arrays (ULAs) consisting of $N_{\text{t}}$ and $N_{\text{r}}$ antennas, respectively, with half-wavelength spacing $\lambda/2$, where $\lambda$ is the wavelength corresponding to the center subcarrier frequency. Self-interference at the BS is assumed to be mitigated by sufficiently isolated TX and RX antenna arrays [40, 41]. The RIS is composed of $N_{\text{s}}$ elements, each with a size of $\xi_{\text{s}}\times\xi_{\text{s}}$. Its phase shifts are denoted by $\bm{\omega}=[\omega_{1},\omega_{2},\ldots,\omega_{N_{\text{s}}}]^{\text{T}}\in\mathbb{C}^{N_{\text{s}}\times 1}$ and dynamically adjusted by the BS in several microseconds [20]. The positions of the TX, RX, and RIS are assumed to be known, whereas other environmental components are unknown but considered static.
@@
The region of interest (ROI) is a predefined area where potential targets may be located. It can be discretized into $N_{\text{v}}$ voxels, and the ROI image is represented as $\bm{\sigma}=[\sigma_{1},\sigma_{2},\ldots,\sigma_{N_{\text{v}}}]^{\text{T}}\in\mathbb{R}^{N_{\text{v}}\times 1}$, where $\sigma_{n_{\text{v}}}$ denotes the scattering coefficient of the $n_{\text{v}}$-th voxel, and $\sigma_{n_{\text{v}}}=0$ if the voxel contains no objects. The voxel side length is denoted by $\xi_{\text{v}}$, which is assumed to be much smaller than $\lambda/2$ [42, 17]. It is assumed that $N_{0}$ symbol intervals are utilized to generate an image. Although multiple subcarriers are available, the frequency index is omitted for simplicity, and the proposed methods can be readily extended to multi-subcarrier scenarios.
(a) User inside the ROI
(b) User outside the ROI
@@
Figure 2: Illustration of the proposed RIS-aided ISAC system for wireless imaging.
II-A Signal and Channel Models for Communication
We assume that the communication user is equipped with a single antenna and is located either inside or outside the ROI. Denote $\mathbf{x}\in\mathbb{C}^{N_{\text{t}}\times 1}$ as the normalized transmitted signal at the TX and $P_{\text{t}}$ as the transmit power. The received signal at the user is given by
${r}_{\text{com}}=\sqrt{P_{\text{t}}}\,\mathbf{h}_{\text{com-}\ast}^{\text{T}}\mathbf{x}+{n}_{\text{com}},$ (2)
where $\mathbf{h}_{\text{com-}\ast}^{\text{T}}\in\mathbb{C}^{1\times N_{\text{t}}}$ represents the multipath channel from the TX to the user, and ${n}_{\text{com}}\in\mathbb{C}$ is the additive Gaussian noise. Here, $\in\{\text{in},\text{out}\}$ denotes the user position, corresponding to the channel for users inside or outside the ROI, respectively.
According to Fig. 2(a), when the communication user is located inside the ROI, it simultaneously serves as the communication terminal and the target being sensed. Thus, $\mathbf{h}_{\text{com-in}}$ is given by
$\mathbf{h}_{\text{com-in}}=\mathbf{h}_{\text{tx-ue}}+\mathbf{h}_{\text{tx-ris-ue}}+\mathbf{h}_{\text{others}},$ (3)
where $\mathbf{h}_{\text{tx-ue}}=g_{\text{com}}\mathbf{h}^{\prime}_{\text{tx-ue}}\in\mathbb{C}^{N_{\text{t}}\times 1}$ denotes the line-of-sight (LOS) path between the TX and the user. Here, $g_{\text{com}}={\lambda\sqrt{G_{\text{com}}}}/{\sqrt{4\pi}}$, and $G_{\text{com}}$ is the combined antenna gain of the TX and the user. The $n_{\text{t}}$-th element in $\mathbf{h}^{\prime}_{\text{tx-ue}}$ is [43]
$h^{\prime}_{\text{tx-ue},n_{\text{t}}}=\frac{e^{-j2\pi\frac{d_{n_{\text{t}},{\text{ue}}}}{\lambda}}}{\sqrt{4\pi}d_{n_{\text{t}},{\text{ue}}}},$ (4)
where $d_{n_{\text{t}},{\text{ue}}}$ denotes the distance between the $n_{\text{t}}$-th TX antenna and the user. The RIS-reflected path is given by
$\mathbf{h}_{\text{tx-ris-ue}}=g_{\text{com}}\mathbf{H}_{\text{tx-ris}}\text{diag}(\widetilde{\bm{\omega}})\mathbf{h}_{\text{ris-ue}},$ (5)
where $\widetilde{\bm{\omega}}=g_{\text{ris}}\bm{\omega}$, and $g_{\text{ris}}=\sqrt{4\pi}\xi_{\text{s}}^{2}/\lambda$ is the scattering gain of the RIS elements [44]. $\mathbf{H}_{\text{tx-ris}}^{\text{T}}$ and $\mathbf{h}_{\text{ris-ue}}^{\text{T}}$ denote the channels from the TX to the RIS and from the RIS to the user, respectively. The elements of both channels follow a similar form to (4). Finally, $\mathbf{h}_{\text{others}}$ includes multipath components involving more than two bounces and other scatterers.
Alternatively, when the user is located outside the ROI, the target inside the ROI introduces additional multipaths into the communication channel, as depicted in Fig. 2(b). In this case, $\mathbf{h}_{\text{com-out}}$ is given by
$\mathbf{h}_{\text{com-out}}=\mathbf{h}_{\text{tx-ue}}+\mathbf{h}_{\text{tx-roi-ue}}+\mathbf{h}_{\text{tx-ris-ue}}\\ +\mathbf{h}_{\text{tx-roi-ris-ue}}+\mathbf{h}_{\text{tx-ris-roi-ue}}+\mathbf{h}_{\text{others}},$ (6)
where the single-bounce path scattered by the ROI is given by
$\mathbf{h}_{\text{tx-roi-ue}}=g_{\text{com}}\mathbf{H}_{\text{tx-roi}}\text{diag}(\bm{\sigma})\mathbf{h}_{\text{roi-ue}},$ (7)
with $\mathbf{H}_{\text{tx-roi}}^{\text{T}}$ and $\mathbf{h}_{\text{roi-ue}}^{\text{T}}$ denoting the channels from the TX to the ROI and from the ROI to the user, respectively. The twice-bounce paths are expressed as
$\mathbf{h}_{\text{tx-roi-ris-ue}}=g_{\text{com}}\mathbf{H}_{\text{tx-roi}}\text{diag}(\bm{\sigma})\mathbf{H}_{\text{roi-ris}}\text{diag}(\widetilde{\bm{\omega}})\mathbf{h}_{\text{ris-ue}},$ (8a)
$\mathbf{h}_{\text{tx-ris-roi-ue}}=g_{\text{com}}\mathbf{H}_{\text{tx-ris}}\text{diag}(\widetilde{\bm{\omega}})\mathbf{H}_{\text{ris-roi}}\text{diag}(\bm{\sigma})\mathbf{h}_{\text{roi-ue}},$ (8b)
where $\mathbf{H}_{\text{roi-ris}}=\mathbf{H}_{\text{ris-roi}}^{\text{T}}$ is the channel between the ROI and the RIS.
II-B Signal and Channel Models for Sensing
@@
The BS performs target sensing by transmitting signals from the TX and receiving the echoes at the RX. We assume that the RIS phases are altered using $K$ different configurations for imaging, given by $\bm{\Omega}=[\bm{\omega}_{1},\bm{\omega}_{2},\ldots,\bm{\omega}_{K}]^{\text{T}}\in\mathbb{C}^{K\times N_{\text{s}}}$. When employing the $k$-th RIS configuration $\bm{\omega}_{k}$ and transmitting the $n_{\text{t}}$-th signal $\mathbf{x}_{n_{\text{t}}}$, the received signal at the RX is expressed as
@@
where $g_{\text{sen}}={\lambda\sqrt{G_{\text{sen}}}}/{\sqrt{4\pi}}$, and $G_{\text{sen}}$ is the combined antenna gain of the TX and the RX. $\mathbf{H}_{\text{tx-roi-ris-rx},k}$ and $\mathbf{H}_{\text{tx-ris-roi-rx},k}$ represent the twice-bounce paths, given by
@@
$\mathbf{H}_{\text{tx-ris-roi-rx},k}$ $=g_{\text{sen}}\mathbf{H}_{\text{tx-ris}}\text{diag}(\widetilde{\bm{\omega}}_{k})\mathbf{H}_{\text{ris-roi}}\text{diag}(\bm{\sigma})\mathbf{H}_{\text{roi-rx}}.$ (12b)
The term ${\mathbf{H}}_{\text{others},k}$ includes multipath components involving more than two bounces or scatterers beyond the ROI and the RIS, representing the unknown portion of ${\mathbf{H}}_{\text{sen},k}$.
The ROI image $\bm{\sigma}$ is expected to be inferred from ${\mathbf{H}}_{\text{sen},k}$, since this channel includes components scattered by the ROI and thus encodes the image information. Moreover, as RIS phases influence the paths scattered by both the ROI and the RIS, altering the RIS configurations provides diverse sensing perspectives toward the target, significantly increasing the effective number of measurements.
II-C ISAC Protocol
The protocol of the ISAC system is depicted in Fig. 3. The TX and the user keep uninterrupted communication by transmitting and receiving downlink communication signals, respectively. During standard communication phases, RIS phases are optimized to enhance communication performance, whereas the RX does not work. Alternatively, during the ISAC interval, the RIS phases are varied across $K$ unique configurations, while the RX receives backscattered echoes, achieving communication-assisted sensing by multiplexing communication signals. While RIS phase optimization for imaging is explored in [17, 9], this study concentrates on the imaging algorithm design, assuming that $\bm{\Omega}$ is randomly generated beforehand. Subsequently, the formulated ROI images can be used to refine communication-oriented RIS phase designs, thereby enhancing subsequent communication performance and realizing sensing-assisted communication, as elaborated in Sec. V.
Figure 3: Illustration of the proposed ISAC protocol.